2.2.14 Your answer is incorrect. Try again Plastic parts produced by an injection-molding operation are checked for conformance to specifications. Each tool contains 13 cavities in which parts are produced, and these parts fall into a conveyor when the press opens. An inspector chooses 2 part(s) from among the 13 at random. Three cavities are affected by a temperature malfunction that results in parts that do not conform to specifications Round your answers to four decimal places a) What is the probability that the inspector finds exactly one nonconforming part? b) What is the probability that the inspector finds at least one nonconforming part?
Answer #1

a)

total = 13

3 are non-confirming

2 are selected

hence

10 confirming and 3 non-confirming

total number to select 2 out of 13 = 13C2 = 78

probability that inspector finds exactly one non-confirming part= 10C1 * 3C1 /78= 30/78 = 0.3846

b)

P(X >= 1) = 1 - P(X = 0)

= 1 - 10C2/13C2

=1-10*9/(13*12)

=0.4231

Please rate